OpenStudy (anonymous):

The graph below is represented by which system of inequalities? -x + 3y < 9 y > -2x + 1 -x + 3y 9 y > -2x + 1 -x + 3y < 9 y -2x + 1 None of these systems represent the graph shown.

OpenStudy (whpalmer4):

Pretend we can't actually see your book, screen, paper, etc.

OpenStudy (jdoe0001):

the graph below looks very blank

OpenStudy (anonymous):

what

OpenStudy (whpalmer4):

there's no graph!

OpenStudy (anonymous):

oh i don't know how to put a graph on here b/c its my first time on this lol

OpenStudy (whpalmer4):

Well, click on the Draw button down at the bottom and draw the graph (including axes)...

OpenStudy (whpalmer4):

where do you see the graph? is it on your screen, in a textbook, on a piece of paper?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

does that work lol

OpenStudy (anonymous):

screen

OpenStudy (whpalmer4):

yes, that worked. well, the first thing I would observe is that you ave a combination of dashed and solid lines. what does that mean?

OpenStudy (anonymous):

this is for the dashed If the problem does not include where it is equal, then you will use a dashed boundary line. In other words, if you have < or >, you will have a dashed line for your boundary line. and solid lines Recall that the graph of a linear equation is a straight line. ... Unless the directions to a problem indicate otherwise, you can use any method to graph it. ... solid line. Dashed boundary line:

OpenStudy (whpalmer4):

bottom line — dashed line and solid line mean different things — one is > or < and the other is \(\ge\) or \(\le\) right?

OpenStudy (whpalmer4):

if you see both in your graph, doesn't that mean your set of inequalities must have both as well?

OpenStudy (anonymous):

oh ok lol sorry

OpenStudy (anonymous):

um yes

OpenStudy (whpalmer4):

and are there any answer choices that fit that description?

OpenStudy (anonymous):

yes

OpenStudy (whpalmer4):

which one(s)?

OpenStudy (anonymous):

all

OpenStudy (whpalmer4):

the ones I see here only have < or >

OpenStudy (anonymous):

oh i thought you mean sorry

OpenStudy (whpalmer4):

so, if we have both dashed and solid lines, we need a system of inequalities that has at least one of \(<,>\) and at least one of \(\le,\ge\) in it. do you have any systems like that in the answer choices?

OpenStudy (anonymous):

yes maybe i don't get that so much lol

OpenStudy (anonymous):

I get the sings but answer choices that have that were i am lost at

OpenStudy (anonymous):

are you going to help me????

OpenStudy (dan815):

lol whpalmer

OpenStudy (dan815):

pretend we cant actually see ur books

OpenStudy (anonymous):

its online test i need help on this

OpenStudy (dan815):

which color do u want?

OpenStudy (anonymous):

blue lol

OpenStudy (dan815):



OpenStudy (dan815):

the dotted line mens its greater and not including that line

OpenStudy (dan815):

the solid line means its including that line and greater

OpenStudy (anonymous):

I see it

OpenStudy (dan815):

first find the equation of both lines

OpenStudy (dan815):